Toy Co Mills is a division of Iowa Woolen Products. For the most recent year, Toy Co had net income of $20,600,000. Included in income was interest expense of $1,490,000. The operation’s tax rate is 40 percent. Total assets of Toy Co Mills are $236,000,000, current liabilities are $52,000,000, and $35,000,000 of the current liabilities are noninterest bearing.

Calculate NOPAT, invested capital, and ROI for Toy Co Mills. (Round ROI to 2 decimal places, e.g. 15.25%.)

NOPAT $

Invested capital $

ROI

%
Answer #1

NOPAT = Net income + Interest

= 20,600,000 + (1,490,000-40%)

= 21,494,000

Invested capital = Total assets - non interest bearing current liabilities

= 236,000,000 - 35,000,000

= 201,000,000

ROI = NOPAT/Invested capital

= 21,494,000/201,000,000

= 10.69%

Comment if you face any issues